A man is now twice as old as his friend. Five years ago, a man was thrice as old as his friend. Determine their present ages

Answer by fractalier(6550)

Call his age now, x.
Cal his friend's age now, y.
Then we have
x = 2y
and five years ago
x-5=3(y-5)
Substitute the first equation into the second and get
2y-5=3(y-5)
2y - 5 = 3y - 15
and
y = 10
so that
x = 20